# Provably Mitigating Overoptimization in RLHF: Your SFT Loss is Implicitly an Adversarial Regularizer

### v1

@@ B.2 Auxiliary Lemmas
$\max_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r)=\min_{r\in\mathcal{R}}\,\max_{\pi\in\Pi}\phi(\pi,r).$ (B.10)
Proof of Lemma B.1.
The foundation of this result is a minimax theorem given by Fan (1953) (Lemma B.2). In our setting, the policy class $\Pi$ is a nonempty set, and the reward model class $\mathcal{R}$ is a nonempty compact Hausdorff space. Furthermore, by our choice of the policy class $\Pi$ in (2.4), $\Pi$ is a convex set. Meanwhile, the function $\phi$ is a concave function of $\pi\in\Pi$ since the dependence on $\pi$ is linear terms plus a negative KL term (concave). Finally, by our assumption, the function $\phi$ is convex-like on the reward model class $\mathcal{R}$ and is also continuous on $\mathcal{R}$. As a result, all the conditions of Lemma B.2 are satisfied and the minimax theorem holds in our problem setup, finishing the proof of Lemma B.1. ∎
Lemma B.2 (Minimax theorem (Fan, 1953)).
@@
$f(x_{3},\cdot)\geq\alpha\cdot f(x_{1},\cdot)+(1-\alpha)\cdot f(x_{2},\cdot)\,\,\text{on $\mathcal{Y}$,}$ (B.11)
@@
$f(\cdot,y_{3})\leq\beta\cdot f(\cdot,y_{1})+(1-\beta)\cdot f(\cdot,y_{2})\,\,\text{on $\mathcal{Y}$.}$ (B.12)
@@
$\max_{x\in\mathcal{X}}\,\min_{y\in\mathcal{Y}}f(x,y)=\min_{y\in\mathcal{Y}}\,\max_{x\in\mathcal{X}}f(x,y).$ (B.13)
Lemma B.3 (Unique maximizer of $\phi$).
@@
$\phi(\pi,r)$ $:=\eta\cdot\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot D_{\mathrm{KL}}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.14)
$\qquad+{\mathcal{L}}_{\mathcal{D}}(r).$ (B.15)
@@
$\max_{\pi\in\Pi}\phi(\pi,r)$ (B.16)
$\qquad=\eta\cdot\max_{\pi\in\Pi}\left\{\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x)}\Big{[}r(x,a^{1})-\beta\cdot D_{\mathrm{KL}}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}\right\}$ (B.17)
$\qquad=\eta\cdot\max_{\pi\in\Pi}\left\{C_{r}-\beta\cdot\mathbb{E}_{x\sim d_{0}}\left[D_{\mathrm{KL}}\left(\pi(\cdot|x)\middle\|\frac{\pi^{\mathrm{ref}}(\cdot|x)\cdot\exp(\beta^{-1}\cdot r(x,\cdot))}{\int_{a^{\prime}\in\mathcal{A}}\mathrm{d}\pi^{\mathrm{ref}}(a^{\prime}|x)\cdot\exp(\beta^{-1}\cdot r(x,a^{\prime}))}\right)\right]\right\},$ (B.18)
@@
$\pi_{r}(\cdot|x)=\frac{\pi^{\mathrm{ref}}(\cdot|x)\cdot\exp(\beta^{-1}\cdot r(x,\cdot))}{\int_{a^{\prime}\in\mathcal{A}}\mathrm{d}\pi^{\mathrm{ref}}(a^{\prime}|x)\cdot\exp(\beta^{-1}\cdot r(x,a^{\prime}))},$ (B.19)
which completes the proof of Lemma B.3. ∎
Appendix C Proofs for Generalization to New Prompt Distributions
@@
$\mathbb{E}_{x\sim d_{1},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{1},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]$ (C.1)
$\qquad=\mathbb{E}_{x\sim d_{1}}\left[\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right]$ (C.2)
$\qquad=\mathbb{E}_{x\sim d_{0}}\left[\frac{d_{1}(x)}{d_{0}(x)}\cdot\left(\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right)\right]$ (C.3)
$\qquad\leq C_{\infty}(d_{0},d_{1})\cdot\mathbb{E}_{x\sim d_{0}}\left[\left|\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right|\right],$ (C.4)
@@
$\mathbb{E}_{x\sim d_{0}}\left[\left|\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right|\right]$ (C.5)
$\qquad=\mathbb{E}_{x\sim d_{0},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{0},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]$ (C.6)
$\qquad=\mathrm{Gap}^{\pi^{\star}}(\pi_{\widehat{r}}),$ (C.7)
Therefore, by combining (C.4) and (C.7), invoking Corollary 5.7, we have that
$\mathbb{E}_{x\sim d_{1},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{1},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\leq\widetilde{\mathcal{O}}\left(\frac{C_{\infty}(d_{0},d_{1})}{\sqrt{N}}\right),$ (C.8)
@@
Appendix D Additional Details on Experiments
D.1 Training Details
@@ D.2 Evaluation Details
D.3 Additional Results on Experiments

### v2

@@ C.2 Auxiliary Lemmas
$\max_{\pi\in\Pi}\,\min_{r\in\mathcal{R}}\phi(\pi,r)=\min_{r\in\mathcal{R}}\,\max_{\pi\in\Pi}\phi(\pi,r).$ (C.10)
Proof of Lemma C.1.
The foundation of this result is a minimax theorem given by Fan (1953) (Lemma C.2). In our setting, the policy class $\Pi$ is a nonempty set, and the reward model class $\mathcal{R}$ is a nonempty compact Hausdorff space. Furthermore, by our choice of the policy class $\Pi$ in (2.4), $\Pi$ is a convex set. Meanwhile, the function $\phi$ is a concave function of $\pi\in\Pi$ since the dependence on $\pi$ is linear terms plus a negative KL term (concave). Finally, by our assumption, the function $\phi$ is convex-like on the reward model class $\mathcal{R}$ and is also continuous on $\mathcal{R}$. As a result, all the conditions of Lemma C.2 are satisfied and the minimax theorem holds in our problem setup, finishing the proof of Lemma C.1. ∎
Lemma C.2 (Minimax theorem (Fan, 1953)).
@@
$f(x_{3},\cdot)\geq\alpha\cdot f(x_{1},\cdot)+(1-\alpha)\cdot f(x_{2},\cdot)\,\,\text{on $\mathcal{Y}$,}$ (C.11)
@@
$f(\cdot,y_{3})\leq\beta\cdot f(\cdot,y_{1})+(1-\beta)\cdot f(\cdot,y_{2})\,\,\text{on $\mathcal{Y}$.}$ (C.12)
@@
$\max_{x\in\mathcal{X}}\,\min_{y\in\mathcal{Y}}f(x,y)=\min_{y\in\mathcal{Y}}\,\max_{x\in\mathcal{X}}f(x,y).$ (C.13)
Lemma C.3 (Unique maximizer of $\phi$).
@@
$\phi(\pi,r)$ $:=\eta\cdot\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot D_{\mathrm{KL}}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (C.14)
$\qquad+{\mathcal{L}}_{\mathcal{D}}(r).$ (C.15)
@@
$\max_{\pi\in\Pi}\phi(\pi,r)$ (C.16)
$\qquad=\eta\cdot\max_{\pi\in\Pi}\left\{\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x)}\Big{[}r(x,a^{1})-\beta\cdot D_{\mathrm{KL}}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}\right\}$ (C.17)
$\qquad=\eta\cdot\max_{\pi\in\Pi}\left\{C_{r}-\beta\cdot\mathbb{E}_{x\sim d_{0}}\left[D_{\mathrm{KL}}\left(\pi(\cdot|x)\middle\|\frac{\pi^{\mathrm{ref}}(\cdot|x)\cdot\exp(\beta^{-1}\cdot r(x,\cdot))}{\int_{a^{\prime}\in\mathcal{A}}\mathrm{d}\pi^{\mathrm{ref}}(a^{\prime}|x)\cdot\exp(\beta^{-1}\cdot r(x,a^{\prime}))}\right)\right]\right\},$ (C.18)
@@
$\pi_{r}(\cdot|x)=\frac{\pi^{\mathrm{ref}}(\cdot|x)\cdot\exp(\beta^{-1}\cdot r(x,\cdot))}{\int_{a^{\prime}\in\mathcal{A}}\mathrm{d}\pi^{\mathrm{ref}}(a^{\prime}|x)\cdot\exp(\beta^{-1}\cdot r(x,a^{\prime}))},$ (C.19)
which completes the proof of Lemma C.3. ∎
Appendix D Proofs for Generalization to New Prompt Distributions
@@
$\mathbb{E}_{x\sim d_{1},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{1},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]$ (D.1)
$\qquad=\mathbb{E}_{x\sim d_{1}}\left[\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right]$ (D.2)
$\qquad=\mathbb{E}_{x\sim d_{0}}\left[\frac{d_{1}(x)}{d_{0}(x)}\cdot\left(\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right)\right]$ (D.3)
$\qquad\leq C_{\infty}(d_{0},d_{1})\cdot\mathbb{E}_{x\sim d_{0}}\left[\left|\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right|\right],$ (D.4)
@@
$\mathbb{E}_{x\sim d_{0}}\left[\left|\mathbb{E}_{a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\right|\right]$ (D.5)
$\qquad=\mathbb{E}_{x\sim d_{0},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{0},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]$ (D.6)
$\qquad=\mathrm{Gap}^{\pi^{\star}}(\pi_{\widehat{r}}),$ (D.7)
Therefore, by combining (D.4) and (D.7), invoking Corollary 5.7, we have that
$\mathbb{E}_{x\sim d_{1},a\sim\pi^{\star}(\cdot|x)}[r^{\star}(x,a)]-\mathbb{E}_{x\sim d_{1},a\sim\pi_{\widehat{r}}(\cdot|x)}[r^{\star}(x,a)]\leq\widetilde{\mathcal{O}}\left(\frac{C_{\infty}(d_{0},d_{1})}{\sqrt{N}}\right),$ (D.8)
@@
Appendix E Additional Details on Experiments
E.1 Training Details
@@ E.2 Evaluation Details
E.3 Additional Results on Experiments
@@
Appendix F Experiments on Math, Reasoning, and Coding Tasks
F.1 Experimental Details
To provide a more comprehensive analysis of the trained LLM, we introduce more benchmarks on the math, reasoning, and coding tasks for evaluations. Specifically, we choose the Grade School Math 8K (GSM8K), AI2 Reasoning Challenge (ARC), and Mostly Basic Python Programming (MBPP) to measure math, reasoning, and coding abilities, respectively. In this section, we focus on the gemma series for the experiments. We do not use chain-of-thought or few shots in all the benchmarks. We compare the greedy decoding result (pass @$1$) on the MBPP benchmark.
Here we use the OpenRLHF codebase (Hu et al., 2024) to implement a new variant of RPO, where the SFT loss regularizer is averaged by the number of tokens of the chosen labels, that is, $(\log\pi_{\theta}(a_{\text{cho}}|x))/|a_{\text{cho}}|$. Such a variant balances the weight of the averaged SFT loss regularizer between the shorter chosen response and the longer one. We set the coefficient for the SFT loss regularizer as $0.2$. We use 8 NVIDIA A100 GPUs for the training and evaluation. The remaining hyperparameters are in Table 7.
Configuration Gemma Series
learning rate 5.0e-7
learning scheduler type cosine with a minimum learning rate
batch size 128
gradient accumulation 8
batch size per device 2
training epoch 2
$\beta$ 0.5
optimizer adamw torch
seed 42
precision bfloat16
Table 7: Training configurations for DPO and RPO for the experiments in Appendix F.
F.2 Experimental Results
Table 8 demonstrates that our proposed method still outperforms or performs equally to the vanilla DPO on these benchmarks of math, reasoning, and coding, which verifies the effectiveness of our proposed method.
Model Name GSM8K ARC MBPP (Pass @$1$)
(%) Easy (%) Challenge (%) Normal (%) Plus (%)
RPO 49.9 79.1 49.8 54.2 46.3
DPO 45.3 75.7 50.0 54.2 43.9
Ref. 45.4 75.0 45.8 50.3 44.2
zephyr-gemma-7b 47.3 77.6 48.6 54.5 44.7
Table 8: The results on GSM8K, ARC, and MBPP. Here, zephyr-gemma-7b is the officially released models trained by DPO and Ref. denotes the reference model zephyr-7b-gemma-sft used for our training. RPO and DPO are trained with the OpenRLHF codebase (Hu et al., 2024) and we average the SFT loss regularizer in RPO by the number of tokens of the chosen response. We do not use chain-of-thought or few shots in all the benchmarks. We compare the greedy decoding result (pass @$1$) for MBPP.